Examples of Applications ton of 1Q scores Ceneral Psychology Testboolk scones tall ithin 15 points of score of 100 distrbution of scores around this mean indicates tha standad deviationo Chaste Apet Chapter 3pdt 4ล. t, the area to the lef, or in. narmul diatribation is 12.wh. hthe area toth, right ofa? t0 lf the area to the right of s in a nannl distrten b0 543,what s the area ta the left of s? prababíity hat 51. Find the prelabty that< 30 52 Fiad the Bo percenca 60P percencle. he probabilty th HOMEWORK
Answer #1

48)
P(X > x) = 1 - P(X < x)
= 1 - 0.123
= 0.877

49)
P(X < x) = 1 - 0.543
= 0.457

For rest of questions

Use
Z = (X - mean)/sd

Then use normal distribution table
to calculate P(Z < z)
Also
P(Z > z) = 1 - P(Z < z)